OpenStudy (i_love_you):

I need help, I'm helping my best friends little sister do some math. But I'm having a problem confirming an answer on this question. Which of the following cereals is the best buy? 1. 28oz. of cereal for $4.29 2. 32oz. of cereal for $4.59 3. 36oz. of cereal for $4.79 4. 24oz. of cereal for $3.59 I'm a little unsure on an answer please help me. Thanks ahead of time! :)

OpenStudy (skullpatrol):

Find the best price per oz.

OpenStudy (skullpatrol):

For example $4.29/28oz

OpenStudy (study312):

4.29 ÷ 28 = 4.59 ÷ 32 = 4.79 ÷ 36 = 3.59 ÷ 24 = can you calculate all of that?

OpenStudy (skullpatrol):

Choose the smallest number of dollars paid per oz.

OpenStudy (i_love_you):

Okay thank you guys. I think I got it now.

OpenStudy (skullpatrol):

Thanks for asking :D

OpenStudy (study312):

I'll give skullpatrol a medal so you can give me a medal

OpenStudy (study312):

np :)